# Flips in colorful triangulations

### v2

@@ 1. Introduction
The associahedron is a polytope of fundamental interest and importance [Lod04, HL07, CSZ15], as it lies at the heart of many recent developments in algebraic combinatorics and discrete geometry; see [PSZ23] and the references therein. In this paper we are specifically interested in its combinatorial structure, namely the graph of its skeleton; see Figure 1. This graph, which we denote by ${\mathcal{G}}_{N}$, has as nodes all triangulations of a convex $N$-gon ($N\geq 3$), and an edge between any two triangulations that differ in a flip operation, which consists of removing an edge shared by two triangles and replacing it by the other diagonal of the resulting 4-gon. The graph ${\mathcal{G}}_{N}$ is isomorphic to the graph that has as nodes all binary trees with $N-2$ vertices, and an edge between any two trees that differ in a tree rotation. Each binary tree arises as the geometric dual of a triangulation, with the root given by ‘looking through’ a fixed outer edge, and flips translate to tree rotations under this bijection; see Figure 2. This dual point of view between triangulations and binary trees, and more generally, between dissections into $(k+1)$-gons and $k$-ary trees, is an essential tool used in our paper.
@@
Another fundamental graph property that we focus on in this paper is Hamiltonicity. To this end, Lucas [Luc87] first proved that ${\mathcal{G}}_{N}$ admits a Hamilton cycle for $N\geq 5$, and a short proof was given by Hurtado and Noy [HN99]. A Hamilton path in ${\mathcal{G}}_{N}$ can be computed efficiently and yields a Gray code ordering of all binary trees by rotations [LRvBR93]. This algorithm is a special case of the more general Hartung-Hoang-Mütze-Williams permutation language framework [HHMW22, HM21, MM23, CMM24].
@@ 1.1. Sagan’s problem and its generalization
For $\alpha=(a,1,1,1,1,1)$, the graph ${\mathcal{F}}_{\alpha}$ has no Hamilton cycle if $a\geq 1$, but a Hamilton path unless $a\in\{1,3\}$.
@@ 1.2. Algorithmic questions and higher arity
We implemented the algorithms mentioned in Theorems 6 and 8 in C++, and made the code available for download, experimentation and visualization on the Combinatorial Object Server [cos].
@@ 1.3. Three colors
Figure 8. Twist operation and the corresponding binary trees.
Figure 9. The flip graphs ${\mathcal{H}}_{6}$ and ${\mathcal{H}}_{9}$.
@@ 3.1. Alternating colors
Figure 12. Flip graphs of colorful triangulations and reduced graphs for $N=4,5,6,7$.
@@
We also note that every quadrangulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{N}$. Indeed, given $R$, then coloring the $N$ points red and blue alternatingly will make all edges colorful. We define a reduced graph ${\mathcal{F}}_{N}^{\prime}$, that has as nodes all quadrangulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{N}^{\prime}$; see Figures 12 and 13. We observe that ${\mathcal{F}}_{N}^{\prime}$ is isomorphic to the flip graph of quadrangulations ${\mathcal{G}}_{N,4}$, i.e., we have ${\mathcal{F}}_{N}^{\prime}\simeq{\mathcal{G}}_{N,4}$. These arguments yield a direct combinatorial proof for the first equality in Theorem 10.
@@ 3.2. General coloring patterns
    Except the edges in $E_{\alpha}$, every monochromatic edge of $T$ is an inner edge.
@@
We refer to a dissection of a convex $N$-gon into $q$ quadrangles and $t$ triangles that contain all the edges of $E_{\alpha}$ as an $\alpha$-angulation, and we write ${\mathcal{D}}_{\alpha}$ for the set of all such dissections. We also note that every $\alpha$-angulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{\alpha}$. Indeed, given $R$, then coloring the $N$ points according to the pattern in (1) will make all edges except the ones in $E_{\alpha}$ colorful. We define a reduced graph ${\mathcal{F}}_{\alpha}^{\prime}$ that has as nodes all $\alpha$-angulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{\alpha}^{\prime}$; see Figures 12, 13 and 15.
@@ 4. Proof of Theorem 5
Figure 18. Illustration of the family of graphs ${\mathcal{F}}_{(a,1,1,1,1,1)}$, with the subgraphs $A$ and $B$ and the two partition classes of $B$ highlighted.
@@
We first argue that ${\mathcal{F}}_{\alpha}$, $a\geq 1$, has no Hamilton cycle. For this observe that the graph has two subgraphs $A$ and $B$, as shown in Figure 18. The subgraph $B$ is bipartite with two partition classes of equal sizes, and it is connected to $A$ only via edges starting from the same partition class. Consequently, if there were a Hamilton cycle in ${\mathcal{F}}_{\alpha}$, then removing the edges between $A$ and $B$ from the cycle would leave a set of spanning paths in $B$ that all start and end in the same partition class of $B$, a contradiction, as the partition classes of $B$ have the same size.
For $a=1$ it is easy to see that there is no Hamilton path; see the top left part of Figure 18 or Figure 12 (the graph ${\mathcal{F}}_{6}$). For $a=2$ a Hamilton path is shown in Figure 12 (in the graph ${\mathcal{F}}_{7}$). For $a=3$ a computer search shows that there is no Hamilton path. For $a\geq 4$, a Hamilton path can be constructed by combining the gadgets shown in Figure 19.
@@ 5.3. The children sequence of a tree
Given a tree $T$ with $n-1$ vertices, in the following we define a sequence $c(T)$ of trees with $n$ vertices, which are obtained by inserting an additional vertex into $T$, specifically into the rightmost branch starting from the root of $T$. This is done in such a way that any two consecutive trees in the sequence $c(T)$ differ in a tree rotation. It turns out that all these rotations are clean, and the sequence $c(T)$ is created by down-rotations of a vertex with the largest label $n$, and the reverse sequence is created by up-rotations of the vertex $n$.
@@
Figure 21. Illustration of subtree sequences and the rightmost branch of a tree ($k=3$).
Given the sequence $\rho(T)=(\rho_{1},\ldots,\rho_{\ell})$, we define $T^{i}:=T_{(\rho_{1},\ldots,\rho_{i})}$ for $i=0,\ldots,\ell$ and $T^{\ell+1}:=T_{\rho(T),1}$. For $i=0,\ldots,\ell$ we have $T^{i}=[S^{i+1},T^{i+1},\varepsilon^{\lambda_{i}}]$ for integers $10\leq\lambda_{i}\leq k-$ 1 if $i<\ell$ and $\lambda_{\ell}=k-1$. Note that $T^{i+1}$ is the rightmost nonempty subtree of $T^{i}$, and $S^{i+1}$ is the sequence of subtrees to the left of it. Also note that $S^{1},\ldots,S^{\ell}$ are nonempty sequences of subtrees, whereas $S^{\ell+1}=\varepsilon$.
We define $T^{i}_{0}:=[T^{i},\varepsilon^{k-1}]$ for all $i=0,\ldots,\ell$, and furthermore, for $j=1,\ldots,\lambda_{i}$ we let $T^{i}_{j}$ be the tree obtained from $T^{i}=[S^{i+1},T^{i+1},\varepsilon^{\lambda_{i}}]$ by replacing the $j$th occurrence of $\varepsilon$ after $T^{i+1}$ by $[\varepsilon^{k}]$.
For $i=0,\ldots,\ell$ and $j=0,\ldots,\lambda_{i}$ we let $c_{i,j}(T)$ be the tree obtained from $T$ by replacing the subtree $T^{i}$ by $T^{i}_{j}$. For $j=0$ this corresponds to inserting a new vertex between the subtree $T^{i}$ and its parent vertex (as root if $i=0$) and making $T^{i}$ the first child of this new vertex. For $j>0$ this corresponds to inserting a new vertex as a child of the root of the subtree $T^{i}$, $j$ positions after the rightmost nonempty subtree.
@@
Combining (3) and (5) shows that any two trees in $C_{i}$ differ in a tree rotation, and if $i<\ell$ then we see from (6) that any two trees in $C_{i}\cup\{c_{i+1,0}(T)\}$ differ in a tree rotation; see Figure 23.
Observe that all these rotations are clean, i.e., if $T$ is a tree with vertex labels $11,\ldots,n-$ 1, then in the tree $c_{i,j}(T)$ the new largest label $n$ is given to the root of the subtree $T^{i}_{0}$ and to the rightmost leaf of $T^{i}_{j}$ for $j=1,\ldots,\lambda_{i}$. I.e., we can think of the trees $c_{i,j}(T)$ as the trees obtained from $T$ by inserting the vertex $n$ in all possible ways on the rightmost branch of $T$ (and the other vertex labels remain unchanged).
@@ 5.4. Rotations between first and last children
Figure 24. Relation between last entries of $c(T)$ and $c(T^{\prime})$ for two trees that differ in a clean tree rotation. Colored edges indicate pairs of corresponding locations of children.
@@ 5.5. Proof of Theorem 7
Figure 26. Hamilton cycle construction by interleaving the children sequences $c(S_{n-1})$ and $c(S_{n-1}^{\prime})$ of two special trees $S_{n-1}$ and $S_{n-1}^{\prime}$. Part (a) shows the construction for $k=2$ (binary trees), and part (b) shows the construction for $k\geq 3$.
@@ 6. Proofs of Theorems 8 and 6
We represent a $k$-ary tree with vertices $1,\ldots,n$ by maintaining an array of length $k$ for the children of each vertex. Furthermore, each vertex has access to its parent vertex. With these data structures, up- and down-rotations by 1 step can be performed in time ${\mathcal{O}}(k)$. The algorithm to generate all trees by rotations is a faithful implementation of the inductive zigzag construction described in Section 5. As we are only computing a Hamilton path in the rotation graph (not a cycle), the interleaving of children sequences for the special trees $S_{n}$ and $S_{n}^{\prime}$ discussed in the proof of Theorem 7 is omitted, which simplifies this computation. Specifically, the selection of the vertex to be rotated next and in which direction can be achieved in time ${\mathcal{O}}(1)$ using simple data structures developed for the Hartung-Hoang-Mütze-Williams permutation language framework; see e.g. [MM23, CMM24].

### v3

@@ 1. Introduction
The associahedron is a polytope of fundamental interest and importance [Lod04, HL07, CSZ15], as it lies at the heart of many recent developments in algebraic combinatorics and discrete geometry; see [PSZ23] and the references therein. In this paper we are specifically interested in its combinatorial structure, namely the graph of its skeleton; see Figure 1. This graph, which we denote by ${\mathcal{G}}_{N}$, has as nodes all triangulations of a convex $N$-gon ($N\geq 3$), and an edge between any two triangulations that differ in a flip operation, which consists of removing an edge shared by two triangles and replacing it by the other diagonal of the resulting 4-gon. The graph ${\mathcal{G}}_{N}$ is isomorphic to the graph that has as nodes all binary trees with $N-2$ vertices, and an edge between any two trees that differ in a tree rotation. Each binary tree arises as the geometric dual of a triangulation, i.e., we place a vertex into every triangle, connect vertices in adjacent triangles by an edge, and the root of the tree is given by ‘looking through’ a fixed outer edge; see Figure 1. Under this bijection between triangulations and binary trees, flips translate to tree rotations; see Figure 2. This dual point of view between triangulations and binary trees, and more generally, between dissections into $(k+1)$-gons and $k$-ary trees, is an essential tool used in our paper.
@@
Another fundamental graph property that we focus on in this paper is Hamiltonicity. To this end, Lucas [Luc87] first proved that ${\mathcal{G}}_{N}$ admits a Hamilton cycle for $N\geq 5$, and a short proof was given by Hurtado and Noy [HN99]. A Hamilton path in ${\mathcal{G}}_{N}$ can be computed efficiently and yields a Gray code ordering of all binary trees by rotations [LRvBR93]. This algorithm is a special case of the more general Hartung-Hoang-Mütze-Williams permutation language framework [HHMW22, HM21, MM23, CMM25].
@@ 1.1. Sagan’s problem and its generalization
For $\alpha=(a,1,1,1,1,1)$, the graph ${\mathcal{F}}_{\alpha}$ has no Hamilton cycle for any $a\geq 1$, but it has a Hamilton path for $a=2$ and any $a\geq 4$.
@@ 1.2. Algorithmic questions and higher arity
Figure 8. Twist operation and the corresponding binary trees.
Figure 9. The flip graphs ${\mathcal{H}}_{6}$ and ${\mathcal{H}}_{9}$.
@@
We implemented the algorithms mentioned in Theorems 6 and 8 in C++, and made the code available for download, experimentation and visualization on the Combinatorial Object Server [cos]. In fact, we also provide a slightly more technical implementation for efficiently computing a Hamilton cycle in ${\mathcal{G}}_{N,k+1}$ for arbitrary $k\geq 2$.
@@ 2. Preliminaries
In this section we introduce some notations that will be used throughout this paper. For the reader’s convenience, some frequently used symbols are summarized in Table 1.
Table 1. Summary of frequently used notations.
Symbol Description
$N$ number of points
$k$ arity of the trees, dissection into $(k+1)$-gons
${\mathcal{G}}_{N}$ associahedron, i.e., flip graph on all triangulations
${\mathcal{T}}_{n,k}$ set of $k$-ary trees with $n$ vertices
${\mathcal{D}}_{N,k+1}$ set of dissections of $N$-gon into $(k+1)$-gons
${\mathcal{G}}_{N,k+1}$ flip graph on dissections ${\mathcal{D}}_{N,k+1}$
${\mathcal{C}}_{N}$ set of colorful triangulations for alternating red-blue coloring
${\mathcal{F}}_{N}\subseteq{\mathcal{G}}_{N}$ flip graph on ${\mathcal{C}}_{N}$
$r(T)$ reduced dissection obtained from triangulation $T$
${\mathcal{F}}_{N}^{\prime}$ flip graph on reduced dissections
$\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ coloring sequence for two colors, $N=\sum_{i=1}^{\ell}\alpha_{i}$
${\mathcal{C}}_{\alpha}$ set of colorful triangulations for coloring sequence $\alpha$
${\mathcal{F}}_{\alpha}$ flip graph on ${\mathcal{C}}_{\alpha}$
$E_{\alpha}$ monochromatic boundary edges
$q$ number of quadrangles
$t$ number of triangles
${\mathcal{D}}_{\alpha}$ set of reduced dissections into quadrangles and triangles
${\mathcal{F}}_{\alpha}^{\prime}$ flip graph on ${\mathcal{D}}_{\alpha}$
${\mathcal{C}}_{N}^{\prime}$ set of colorful triangulations for alternating red-blue-green coloring
${\mathcal{H}}_{N}$ flip graph on ${\mathcal{C}}_{N}^{\prime}$
$\varepsilon$ empty string
$\operatorname{rev}(x)$ reverse of sequence $x$
$f(x)$, $\ell(x)$ first and last entry of sequence $x$
$v(G)$, $\Delta(G)$ number of vertices of graph $G$, maximum degree of $G$
$G\simeq H$ isomorphic graphs $G$ and $H$
$Q_{d}$ $d$-dimensional hypercube
$\nu_{T}(S)$ subtree sequence for subtree $S$ of $T$
$\rho(T)$ rightmost branch of $T$
$T\sim T^{\prime}$ two trees that differ in a rotation
$c(T)$ children sequence of tree $T$
@@ 3.1. Alternating colors
Figure 12. Flip graphs of colorful triangulations and reduced graphs for $N=4,5,6,7$. A Hamilton path in ${\mathcal{F}}_{7}$ is highlighted.
@@
We also note that every quadrangulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{N}$. Indeed, given $R$, then coloring the $N$ points red and blue alternatingly will make all edges colorful. We define a reduced graph ${\mathcal{F}}_{N}^{\prime}$ that has as nodes all quadrangulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{N}^{\prime}$; see Figures 12 and 13. We observe that ${\mathcal{F}}_{N}^{\prime}$ is isomorphic to the flip graph of quadrangulations ${\mathcal{G}}_{N,4}$, i.e., we have ${\mathcal{F}}_{N}^{\prime}\simeq{\mathcal{G}}_{N,4}$. These arguments yield a direct combinatorial proof for the first equality in Theorem 10.
@@ 3.2. General coloring patterns
    Apart from the edges in $E_{\alpha}$, every monochromatic edge of $T$ is an inner edge.
@@
We refer to a dissection of a convex $N$-gon into $q$ quadrangles and $t$ triangles that contain all the edges of $E_{\alpha}$ as an $\alpha$-angulation, and we write ${\mathcal{D}}_{\alpha}$ for the set of all such dissections. We also note that every $\alpha$-angulation $R$ on $N$ points equals $r(T)$ for some colorful triangulation $T\in{\mathcal{C}}_{\alpha}$. Indeed, given $R$, then coloring the $N$ points according to the pattern in (1) will make all edges apart from the ones in $E_{\alpha}$ colorful. We define a reduced graph ${\mathcal{F}}_{\alpha}^{\prime}$ that has as nodes all $\alpha$-angulations on $N$ points, and for any two colorful triangulations $T$ and $T^{\prime}$ that differ in a flip of a colorful edge, we add an edge between $r(T)$ and $r(T^{\prime})$ in ${\mathcal{F}}_{\alpha}^{\prime}$; see Figures 12, 13 and 15.
@@ 4. Proof of Theorem 5
Figure 18. Top: Illustration of the family of graphs ${\mathcal{F}}_{(a,1,1,1,1,1)}$, with the subgraphs $A$ and $B$ and the two partition classes of $B$ highlighted. Bottom: The reduced graph ${\mathcal{F}}_{(a,1,1,1,1,1)}^{\prime}$ for $a=4$, explaining the general structure.
@@
We first argue that ${\mathcal{F}}_{\alpha}$, $a\geq 1$, has no Hamilton cycle. For this observe that the graph has two subgraphs $A$ and $B$, as shown in Figure 18. Specifically, the subgraph $A$ is induced by all triangulations $T\in{\mathcal{C}}_{\alpha}$ in which the point $a+3$ has degree 2 in the dissection $r(T)$ (i.e., it only has the previous point $a+2$ and the next point $a+4$ as neighbors). The subgraph $B$ is induced by all other triangulations from ${\mathcal{C}}_{\alpha}$, and the structure and interaction of both subgraphs can be deduced from the bottom part of the figure. The subgraph $B$ is bipartite with two partition classes of equal sizes, and it is connected to $A$ only via edges starting from the same partition class. Consequently, if there were a Hamilton cycle in ${\mathcal{F}}_{\alpha}$, then removing the edges between $A$ and $B$ from the cycle would leave a set of spanning paths in $B$ that all start and end in the same partition class of $B$, a contradiction, as the partition classes of $B$ have the same size.
For $a=1$ it is easy to see that there is no Hamilton path; see the top left part of Figure 18 or Figure 12 (the graph ${\mathcal{F}}_{6}$). For $a=2$ a Hamilton path is shown in Figure 12 (in the graph ${\mathcal{F}}_{7}$). For $a=3$ a computer search shows that there is no Hamilton path. For $a\geq 4$, a Hamilton path can be constructed by combining the gadgets shown in Figure 19. ∎
@@ 5.3. The children sequence of a tree
Given a tree $T$ with $n-1$ vertices, in the following we define a sequence $c(T)$ of trees with $n$ vertices, which are obtained by inserting an additional vertex into $T$ at diffent positions of the rightmost branch starting from the root. This is done in such a way that any two consecutive trees in the sequence $c(T)$ differ in a tree rotation. It turns out that all these rotations are clean, and the sequence $c(T)$ is created by down-rotations of a vertex with the largest label $n$, and the reverse sequence is created by up-rotations of the vertex $n$; see Figure 22 below.
@@
Figure 21. Illustration of subtree sequences and the rightmost branch of a tree ($k=3$). The sequence of all children created from $T$ is displayed in Figure 22.
Given the sequence $\rho(T)=(\rho_{1},\ldots,\rho_{\ell})$, we define a sequence of subtrees of $T$ along its rightmost branch by $T^{i}:=T_{(\rho_{1},\ldots,\rho_{i})}$ for $i=0,\ldots,\ell$. Furthermore, we define $T^{\ell+1}:=T_{\rho(T),1}$, which is the leftmost subtree of $T_{\rho(T)}$. For $i=0,\ldots,\ell$ we have $T^{i}=[S^{i+1},T^{i+1},\varepsilon^{\lambda_{i}}]$ for integers $\lambda_{i}$ and sequences of subtree $S^{i+1}$ that are determined by $T$. Specifically, $T^{i+1}$ is the rightmost nonempty subtree of $T^{i}$, $\lambda_{i}$ is the number of empty subtrees to the right of it, and $S^{i+1}$ is the sequence of subtrees to the left of it. Note that $S^{1},\ldots,S^{\ell}$ are nonempty sequences of subtrees, whereas $S^{\ell+1}=\varepsilon$, i.e., we have $10\leq\lambda_{i}<k-$ 1 if $i<\ell$ and $\lambda_{\ell}=k-1$.
We define $T^{i}_{0}:=[T^{i},\varepsilon^{k-1}]$ for all $i=0,\ldots,\ell$, i.e., $T^{i}_{0}$ is obtained from $T^{i}$ by adding a new root vertex and making $T^{i}$ its leftmost child. Furthermore, for $j=1,\ldots,\lambda_{i}$ we let $T^{i}_{j}$ be the tree obtained from $T^{i}=[S^{i+1},T^{i+1},\varepsilon^{\lambda_{i}}]$ by replacing the $j$th occurrence of $\varepsilon$ after $T^{i+1}$ by $[\varepsilon^{k}]$. In words, $T^{i}_{j}$ is obtained from $T^{i}$ by inserting a new childless vertex $j$ positions to the right of the subtree $T^{i+1}$.
For $i=0,\ldots,\ell$ and $j=0,\ldots,\lambda_{i}$ we let $c_{i,j}(T)$ be the tree obtained from $T$ by replacing the subtree $T^{i}$ by $T^{i}_{j}$. For $j=0$ this corresponds to inserting a new vertex between the subtree $T^{i}$ and its parent vertex (as root if $i=0$) and making $T^{i}$ the first child of this new vertex. For $j>0$ this corresponds to inserting a new vertex as a child of the root of the subtree $T^{i}$, $j$ positions to the right of the rightmost nonempty subtree.
@@
Combining (3) and (5) shows that any two trees in $C_{i}$ differ in a tree rotation, and if $i<\ell$ then we see from (6) that any two trees in $C_{i}\cup\{c_{i+1,0}(T)\}$ differ in a tree rotation; see Figure 23. Observe that all these rotations are clean, i.e., if $T$ is a tree with vertex labels $11,\ldots,n-$ 1, then in the tree $c_{i,j}(T)$ the new largest label $n$ is given to the root of the subtree $T^{i}_{0}$ and to the rightmost leaf of $T^{i}_{j}$ for $j=1,\ldots,\lambda_{i}$. I.e., we can think of the trees $c_{i,j}(T)$ as the trees obtained from $T$ by inserting the vertex $n$ in all possible ways on the rightmost branch of $T$ (and the other vertex labels remain unchanged).
@@ 5.4. Rotations between first and last children
Figure 24. Relation between last entries of $c(T)$ and $c(T^{\prime})$ for two trees that differ in a clean tree rotation. Edges of the same color lead to the same subtrees. The arrows indicate down-rotations of the vertex $b$ by 1 step.
@@ 5.5. Proof of Theorem 7
Figure 26. Illustration of the construction of a Hamilton cycle in the proof of Theorem 7. Specifically, the figure shows the interleaving of the children sequences $c(S_{n-1})$ and $c(S_{n-1}^{\prime})$ of two special trees $S_{n-1}$ and $S_{n-1}^{\prime}$. As in Figure 25, the children sequences are displayed below each tree from top to bottom (corresponding to forward order of the sequence). Part (a) shows the construction for $k=2$ (binary trees), and part (b) shows the construction for $k\geq 3$, subdivided into the two cases $n\geq 4$ (top right) and $n=3$ (bottom right).
@@ 6. Proofs of Theorems 8 and 6
We represent a $k$-ary tree with vertices $1,\ldots,n$ by maintaining an array of length $k$ for the children of each vertex. Furthermore, each vertex has access to its parent vertex. With these data structures, up- and down-rotations by 1 step can be performed in time ${\mathcal{O}}(k)$. The algorithm to generate all trees by rotations is a faithful implementation of the inductive zigzag construction described in Section 5. As we are only computing a Hamilton path in the rotation graph (not a cycle), the interleaving of children sequences for the special trees $S_{n}$ and $S_{n}^{\prime}$ discussed in the proof of Theorem 7 is omitted, which simplifies this computation. Specifically, the selection of the vertex to be rotated next and in which direction can be achieved in time ${\mathcal{O}}(1)$ using simple data structures developed for the Hartung-Hoang-Mütze-Williams permutation language framework; see e.g. [MM23, CMM25].